OpenStudy (anonymous):

May I have some help? Out of the 150 students at a summer camp, 72 signed up for canoeing. There were 23 students who signed up for trekking and 13 of those students also signed up for canoeing. Use a two-way table to organize the information and answer the following question: Approximately what percentage of students signed up for neither canoeing nor trekking?

OpenStudy (anonymous):

I just need someone to help me through the process, it's a little difficult.

OpenStudy (anonymous):

Here are the choices: 45% 68% 28% 42%

OpenStudy (anonymous):

@niksva what do you think the answer is? I say the answer's 42%.

OpenStudy (anonymous):

@Loser66 May I please have some help?

OpenStudy (anonymous):

@ranga May I have some help?

OpenStudy (anonymous):

how did u get 42%?

OpenStudy (anonymous):

I added 72 + 13 + 23 = 108, and then I added 42 to 108 and got 150.

OpenStudy (anonymous):

@phi may I have some help?

OpenStudy (anonymous):

May I have some help @Agent47

OpenStudy (anonymous):

the method which i m using on the basis of P(T union C) = P(T) + P(C) - P(T intersection C) where T is trekking and C is canoeing

OpenStudy (anonymous):

So, it's not 42. What's the answer you got then? :-)

OpenStudy (anonymous):

after using this method i m getting P( T union C) = 46% in order to determine people not signed up for canoeing and trekking P = (100- 46)% = 54%

OpenStudy (anonymous):

Sadly, that's not one of the choices. :(

OpenStudy (anonymous):

yay thats not in the choices

OpenStudy (anonymous):

So, what would you get for your answer out of the 4 choices: 45% 68% 28% 42%

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i m getting 54%

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

The Joint Relative Frequency = the proportional values in the interior of the table. JRF = Joint frequency __________________ total size of sample 13 ___ 150 ???

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

42 over 150 and you get 0.28, move the decimal over 2 places, and you get 28%.

OpenStudy (anonymous):

I got 42 by: 150 - 72 - 23- 13 = 42

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

how do you keep getting 45

OpenStudy (anonymous):

in this ques it has been said 13 of those students so it means that out of 23 students , 13 students signed up for canoeing too m i interpreting this right or not?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

then i m getting 54 %

OpenStudy (anonymous):

Did you divide the answer by 150?

OpenStudy (anonymous):

yup i did

OpenStudy (anonymous):

so what was it? @niksva @grama-pimpsta

OpenStudy (secret-ninja):

The numbers 72 23 and 13 are just there to confuse you. If you add them all together, you get 108. You now just have to subtract 108 from 150. Now you get 42, and 42% is one of the choices, but it is also there to trick you. You have the number, not the percent. You just have to find what percent is 42 out of 150. That would be 28% or C. :)

OpenStudy (secret-ninja):

@Grama-Pimpsta @niksva @zion_15